# Electronic structure, spin-orbit interaction and electron–phonon coupling of triangular adatom lattices on semiconductor substrates

### v1

@@ I Introduction
The material class of a 1/3 monolayer of group-IV adsorbates (Pb, Sn) on semiconductor substrates has a long history. The 1/3 monolayer coverage leads to a $\sqrt{3}\times\sqrt{3}R30^{\circ}$ reconstruction where the atoms sit on the $T_{4}$ sites in a triangular lattice. Thus three out of their four valence bonds are saturated with the substrate, while the fourth orbital remains “dangling” and leads to a half-filled surface band, susceptible to electron-electron interactions. The symmetry-broken low-temperature ground states have fascinated researchers for decades. It started with Pb/Ge(111) and the observation of a low-temperature surface charge density wave Carpinelli et al. (1996). The same group of researchers found a similar phenomenology for the sister compound Sn/Ge(111) Carpinelli et al. (1997). However, the transition from the room temperature $\sqrt{3}\times\sqrt{3}$ phase to the low temperature $3\times 3$ phase associated with charge-ordering could not be explained using first-principle methods, in contrast to Pb/Ge(111) Carpinelli et al. (1996). The authors conjectured Carpinelli et al. (1997) that electron correlations might be responsible for the charge ordering in Sn/Ge(111). Also the role of defects was discussed in this context Weitering et al. (1958); Melechko et al. (2000).
@@
The importance of electron-electron interactions was discussed from the beginning Carpinelli et al. (1997); Santoro et al. (1999), but probably the first paper to propose and explicate a Mott–Hubbard scenario was Ref. Profeta and Tosatti, 2007. By performing local density and Hubbard-$U$ approximation, they proposed both Sn/Ge(111) and Sn/Si(111) as Mott insulating materials. Ref. Cortés et al., 2013 also reported the charge-ordered $3\times 3$ phase in Sn/Ge(111), and characterized it as an intermediate phase between the metallic $3\times 3$ and Mott insulator phases. The importance of spin-orbit coupling (SOC) was emphasized in Ref. Tresca and Calandra, 2021 for Pb/Ge(111) and Pb/Si(111).
@@
Sn/Si(111) was originally believed to feature surface-metallicity, but micro-four-point-probe conductivity measurements revealed insulating behavior from room temperature all the way down to low temperatures Hirahara et al. (2009); Xiong et al. (2022). Doping by partially replacing Sn atoms with In or Na deposition, leads rather suddenly to a metallic conductivity suggesting a small energy gap of the undoped Sn/Si(111) system. Angle-resolved photoemission spectroscopy performed on Sn/Si(111) revealed that the measured spectral function is incompatible with both $3\times 3$ and $\sqrt{3}\times\sqrt{3}$ reconstruction, but compatible with a $2\sqrt{3}\times 2\sqrt{3}$ phase associated with a row-wise antiferromagnetic order Li et al. (2013). The accompanying many-body calculations estimated a quite sizeable Hubbard-$U=0.66$eV. The insulating ground state of Sn/Si(111) was also suggested as a consequence of a Slater-type insulator via band magnetism Lee et al. (2014). Ref. Hansmann et al., 2013 performed a fully first-principle approach, predicting the size of local and non-local Coulomb interactions and derived phase diagrams for various Y/Si(111) compounds (Y=Si, C, Sn, Pb). For Sn/Si(111) they found local Hubbard-$U=1$eV and nearest-neighbor $V_{1}=0.5$eV compared to a bandwidth $W\sim 0.5$eV, further establishing Sn/Si(111) as a Mott insulator. The importance of SOC for Y/Si(111) was emphasized in Ref. Badrtdinov et al., 2016, where SOC combined with strong electron correlations was even proposed to stabilize magnetic skyrmions due to Dzyaloshinskii-Moriya interactions.
@@
Figure 1: DFT band structure and extracted tight binding model for a Sn/Si(111), b Pb/Si(111), c Sn/SiC(0001) and d Pb/SiC(0001). DFT+$U$ has been used to match experimental band gaps for the substrates. Tight binding models have been constructed using the hopping parameters listed in Tab. 1 and SOC splitting in Tab. 2, 3 and 4.
Successful hole-doping of Sn/Si(111) Ming et al. (2017) and indications for superconductivity Ming et al. (2018) were reported, supporting the scenario of a doped Mott insulator. Shortly after, the discovery of superconductivity with $T_{c}=4.7$K in boron-doped Sn/Si(111) marks a breakthrough Wu et al. (2020). Follow-up theory work suggested that the superconducting ground state might be unconventional and constitute chiral topological superconductivity Wolf et al. (2022); Biderang et al. (2022), its nature depending on filling, local and non-local Hubbard interactions. Most recently, experiments on B-doped Sn/Si(111) reported the realization of chiral superconductivity Ming et al. (2023) with critical temperatures instead close to $T_{c}\sim 10$K.
@@
The aim of this paper is to revisit the band structures of the materials Sn/Si(111), Pb/Si(111) and Sn/SiC(0001) in their $\sqrt{3}\times\sqrt{3}$-reconstructued surface phase with a focus on the strength of spin-orbit coupling (SOC) in Sec. II; we complement our investigation with Pb/SiC(0001) whose $\sqrt{3}\times\sqrt{3}$-phase has not been reported so far. We do not investigate many-body ground states involving magnetic or charge order using DFT, but rather study and analyze the non-interacting metallic bandstructures. We further analyze phonon spectra of these materials in Sec. III in order to evaluate the electron-phonon coupling (EPC) strength $\lambda$ in Sec. IV A relativistic study of EPC for Pb/Si(111) is presented in Sec. V. The discussion of our results is presented in Sec. VI, before the end with an outlook in Sec. VII.
@@ II Results I: Electronic structure
The electronic band structure of the four materials with the exception of Pb/SiC(0001) has been reported previously Tresca et al. (2018); Glass et al. (2015); Badrtdinov et al. (2016, 2018); Hansmann et al. (2013); Adler et al. (2019). Here we revisit and compare the four materials. We are interested in their metallic, non-interacting surface bandstructures. By deriving precise tight-binding models of these bands, we provide valuable input for any future many-body analysis of magnetic, charge-ordered, superconducting or otherwise ordered ground states. While the DFT bandstructures of the four materials are rather similar, it turns out that they realize different regimes of SOC ranging from 10% to 45% of the nearest-neighbor hopping amplitude $t_{1}$. This might have major implications for the (strongly) correlated many-body ground states of these materials, as pointed out in a recent study for the plain-vanilla triangular-lattice Rashba-Hubbard model Bunney et al. .
@@
Table 2: Rashba SOC parameters using DFT+$U$. Parameters are given as ratio of $\alpha_{n}/t_{1}$ for $n$th nearest neighbour splitting.
@@
We use the ab initio package VASP Kresse and Furthmüller (1996) within the Perdew-Burke-Ernzerhof (PBE) generalised gradient approximation (GGA) Perdew et al. (1996) to compute the relativistic electronic bandstructures of the four materials. In particular, we employ DFT$+U$ using the approach of Dudarev et. al. Dudarev et al. (1998) where the Coulomb potential $U$ is applied only to the Si or SiC substrate atoms, but not on the Sn and Pb adatoms. The main effect of $U$ is to vary the gap size of the semiconductor substrate foo ; only through changes of the gap size is the metallic surface band influenced. In these calculations we have used a plane-wave cutoff energy of 600 eV and energy convergence criteria of $10^{-8}$eV. For X/SiC(111) $7\times 7\times 1$ $\Gamma$-centered grids were used to sample the Brillouin zone, whereas for X/Si(111) $10\times 10\times 1$ $\Gamma$-centered grids were used for better tight-binding fits close to the $\Gamma$-point. The DFT$+U$ bandstructures along with the extracted tight-binding bands (discussed in detail below) are shown in Fig. 1, zoomed in on the energy range of the metallic surface band. We find the bandwidths $W/t_{1}=9.71$ ($W=0.51$eV) for Sn/Si(111), $W/t_{1}=7.69$ ($W=0.53$eV) for Pb/Si(111), $W/t_{1}=9.29$ ($W=0.40$eV) for Sn/SiC(0001) and $W/t_{1}=9.97$ ($W=0.40$eV) for Pb/SiC(0001). In Appendix A.1 we show in Fig. 7 the same bandstructure plots with larger energy range involving substrate bands.
@@
Table 3: Radial Rashba SOC parameters present for 4th-, 7th-, 9th-nearest neighbour shells.
The four considered materials constitute a case where the metallic surface band is energetically completely separated from the substrate bands, and the description as a single-band tight-binding model is fully justified. The corresponding tight-binding parameters are listed in Tab. 1 and in Tab. 2.
@@
Table 4: Valley Zeeman parameters present for nearest neighbour shells with sites lying off the $C_{3v}$ mirror planes determined by the bulk substrate.
While the isolated triangular lattice of surface adatoms carries a $D_{6h}$ point group symmetry, the presence of the Si(111) or SiC(0001) substrate breaks inversion symmetry, and reduces the rotational symmetry from $C_{6}$ to $C_{3}$. This reduces the point group symmetry to $C_{3v}$. Group theoretic techniques were used to decompose the Hamiltonian terms, implementational details of which we delegate to Appendix A.1. Further symmetry analysis of the model has consequences discussed further below. Interestingly, we find for the nearest-neighbor Rashba amplitude $\alpha_{1}/t_{1}$ the value 0.10 for Sn/Si(111), 0.17 for Sn/SiC(0001), 0.29 for Pb/Si(111) and 0.44 for Pb/SiC(0001). The SOC splitting in Sn/SiC(0001) was previously reported to be $13.7\%$Glass et al. (2015). We attribute the increase in SOC strength in the X/SiC(0001) materials compared to their X/Si(111) counterparts due to the lack of inversion symmetry in the bulk SiC compared to bulk Si. The tight-binding Hamiltonian for our systems can be written as
@@
We only find finite $\phi$ mixings for further-nearest neighbour shells containing 12 sites, where the radial strengths $\beta_{ij}$ shown in Tab. 3 are comparable to the conventional $\alpha_{ij}$. We also find what appears to be a valley Zeeman term, given by $H_{\rm{VZ}}$. This term is finite for nearest neighbours that do not lie on any of the mirror planes of the lattice. The sign of the term is governed by the form factor $\nu_{ij}=-{\rm{sgn}}(\sin(3\theta_{ij}))$. The strengths $J_{ij}$ given in Tab. 4 show that this effect is quite weak, with $J_{2}/t_{1}$ less than 1% in all materials.
@@
Figure 2: Top-down view of Sn/Si(111) with nearest neighbour Sn surface atoms (silver). Surface Si atoms (blue) have been darkened, emphasising how the bulk determines the choice of mirror planes (black lines) of the $C_{3v}$ symmetry.
@@
Some of our derived tight-binding parameter deviate significantly from previously reported ones Li et al. (2013); Adler et al. (2019); Badrtdinov et al. (2016, 2018); Glass et al. (2015); Hansmann et al. (2013). Most of the change is caused by the inclusion of $U$ in the DFT calculations. By increasing the gap size in order to match the experimentally observed values, the propagation of the Wannier functions is suppressed. The larger band gap acts as a wall resulting in highly localised Wannier functions and a decrease in hybridisation with the bulk. Consequently our Wannier functions (shown in Fig. 3 a-d) are much more compact as compared with previous studies Badrtdinov et al. (2016); Glass et al. (2015). This leads to a reduction in the longer-ranged hoppings, e.g.,  $t_{2}/t_{1}$ from $-0.389$Adler et al. (2019) to $-0.272$ for Sn/Si(111), corresponding to a 30% decrease. While second and further neighbor hoppings are still non-negligible, these materials are closer to the nearest-neighbor triangular lattice than previously thought. More importantly, Rashba SOC can be quite sizeable for some of these materials. We note, however, that also our DFT calculations without inclusion of any $U$ led to some discrepancies with the literature. We assume that this is due to the higher resolution used in this work, as well as differing exchange-correlation functionals in the DFT calculations. In Appendix A.3 we compare in Tab. 6 our derived parameters with previous works.
@@
As previously mentioned, the presence of the substrate reduces the point group symmetry of the surface triangular lattice from $D_{6h}$ to $C_{3v}$, by breaking inversion symmetry and reducing the $C_{6}$ rotation symmetry to a $C_{3}$. This can be seen in Fig. 2, which shows that the substrate also preserves the mirror planes which do not/do go through the bonds. We can further understand the terms in the Hamiltonian (1) in three subsets by a hierarchy of symmetry breaking. The kinetic Hamiltonian $H_{0}$ preserves the full $D_{6h}$ point symmetry group. The Rashba SOC terms $H_{R}$ only break inversion symmetry and therefore have a $C_{6v}$ point group symmetry. Note that the radial Rashba effect is only $C_{6v}$ symmetric on 12 site coordination shells. The valley Zeeman-type terms $H_{\rm{VZ}}$ additionally break the the rotational symmetry and have a $C_{3v}$ symmetry. Anisotropic valley Zeeman terms symmetry under $C_{3v}$ are only possible on nearest-neighbor shells where the mirror planes do not intersect with bonds, and hence are only finite on these shells. We also note that additional Rashba coupling terms would be allowed under the reduction to a $C_{3v}$ point group symmetry, but these were not observed.
@@
This symmetry hierarchy serves as one explanation of the relative strengths of the fitted parameters Winkler (2003). Initially employed to understand bandstructures of bulk semiconductors under external magnetic field or uniaxial strain Trebin et al. (1979); Suzuki and Hensel (1974), the symmetry hierarchy is the principle that terms in the Hamiltonian with lower symmetry typically have smaller prefactors than those with higher symmetry. While this would generally mean that the Hamiltonian can be well approximated by the higher symmetry terms alone, the lower symmetry terms may still lead to important physical effects. In this case, the Rashba and the anisotropic valley Zeeman type terms are small, but follow Moriya’s rules for antisymmetric anisotropic spin exchange Moriya (1960a). The Dzyaloshinskii–Moriya interaction Moriya (1960a, b); Dzyaloshinsky (1958) is an important factor for explaning weak anisotropic ferromagnetism in bulk materials, and is shown to drive a first order phase transition to a weakly ferromagnetic phase above a (second-order transition to) an antiferromagnetic phase. This types of interactions remains unexplored in the realm of surface states, and could have potential consequences for known spin-density wave phases in these materials Li et al. (2013).
@@
Figure 3: Maximally localised Wannier functions (MLWFs) for a Sn/Si(111), b Pb/Si(111), c Sn/SiC(0001) and d Pb/SiC(0001) are shown from top and side views. Adatoms are shown in grey, Sn in blue, C in brown, and H in pink. MLWFs are calculated using DFT+$U$ without the effects of SOC. We see low hybridisation with the bulk in all materials, where the MLWFs only penetrate past the first layer.

### v2

@@ I Introduction
The material class of a 1/3 monolayer of group-IV adsorbates (Pb, Sn) on semiconductor substrates has a long history. The 1/3 monolayer coverage leads to a $\sqrt{3}\times\sqrt{3}R30^{\circ}$ reconstruction where the atoms sit on the $T_{4}$ sites in a triangular lattice. Thus three out of their four valence bonds are saturated with the substrate, while the fourth orbital remains “dangling” and leads to a half-filled surface band, susceptible to electron-electron interactions. The symmetry-broken low-temperature ground states have fascinated researchers for decades. It started with Pb/Ge(111) and the observation of a low-temperature surface charge density wave Carpinelli et al. (1996). The same group of researchers found a similar phenomenology for the sister compound Sn/Ge(111) Carpinelli et al. (1997). However, the transition from the room temperature $\sqrt{3}\times\sqrt{3}$ phase to the low temperature $3\times 3$ phase associated with charge-ordering could not be explained using first-principles methods, in contrast to Pb/Ge(111) Carpinelli et al. (1996). The authors conjectured Carpinelli et al. (1997) that electron correlations might be responsible for the charge ordering in Sn/Ge(111). Also the role of defects was discussed in this context Weitering et al. (1999); Melechko et al. (2000).
@@
The importance of electron-electron interactions was discussed from the beginning Carpinelli et al. (1997); Santoro et al. (1999), but probably the first paper to propose and explicate a Mott–Hubbard scenario was Ref. Profeta and Tosatti, 2007. By performing local density and Hubbard-$U$ approximation, they proposed both Sn/Ge(111) and Sn/Si(111) as Mott insulating materials. In Ref. Schuwalow et al., 2010 electronic correlations in Sn/Ge(111) and Sn/Si(111) were considered beyond perturbative approximations within dynamical mean-field theory. Ref. Cortés et al., 2013 also reported the charge-ordered $3\times 3$ phase in Sn/Ge(111), and characterized it as an intermediate phase between the metallic $3\times 3$ and Mott insulator phases. The importance of spin-orbit coupling (SOC) was emphasized in Ref. Tresca and Calandra, 2021 for Pb/Ge(111) and Pb/Si(111).
@@
Sn/Si(111) was originally believed to feature surface-metallicity, but micro-four-point-probe conductivity measurements revealed insulating behavior from room temperature all the way down to low temperatures Hirahara et al. (2009); Xiong et al. (2022). Doping by partially replacing Sn atoms with In or Na deposition, leads rather suddenly to a metallic conductivity suggesting a small energy gap of the undoped Sn/Si(111) system. Angle-resolved photoemission spectroscopy performed on Sn/Si(111) revealed that the measured spectral function is incompatible with both $3\times 3$ and $\sqrt{3}\times\sqrt{3}$ reconstruction, but compatible with a $2\sqrt{3}\times 2\sqrt{3}$ phase associated with a row-wise antiferromagnetic order Li et al. (2013). The accompanying many-body calculations estimated a quite sizeable Hubbard-$U=0.66$eV. The insulating ground state of Sn/Si(111) was also suggested as a consequence of a Slater-type insulator via band magnetism Lee et al. (2014). Ref. Hansmann et al., 2013 performed a fully first-principles approach, predicting the size of local and non-local Coulomb interactions and derived phase diagrams for various Y/Si(111) compounds (Y=Si, C, Sn, Pb). For Sn/Si(111) they found local Hubbard-$U=1$eV and nearest-neighbor $V_{1}=0.5$eV compared to a bandwidth $W\sim 0.5$eV, further establishing Sn/Si(111) as a Mott insulator. The importance of SOC for Y/Si(111) was emphasized in Ref. Badrtdinov et al., 2016, where SOC combined with strong electron correlations was even proposed to stabilize magnetic skyrmions due to Dzyaloshinskii-Moriya interactions.
@@
Figure 1: DFT band structure and extracted tight binding model for a Sn/Si(111), b Pb/Si(111), c Sn/SiC(0001) and d Pb/SiC(0001). DFT+$U$ has been used to match experimental band gaps for the substrates. Tight binding models have been constructed using the hopping parameters listed in Tab. 1 and SOC splitting in Tab. 2, 3 and 4. Fermi surfaces at half-filling calculated from the tight-binding models are included as insets.
Successful hole-doping of Sn/Si(111) Ming et al. (2017) and indications for superconductivity Ming et al. (2018) were reported, supporting the scenario of a doped Mott insulator. At the same time, theoretical work proposed chiral $d$-wave superconductivity in Sn/Si(111) Cao et al. (2018). Shortly after, the discovery of superconductivity with $T_{c}=4.7$K in boron-doped Sn/Si(111) marks a breakthrough Wu et al. (2020). Follow-up theory work suggested that the superconducting ground state might be unconventional and constitute chiral topological superconductivity Wolf et al. (2022); Biderang et al. (2022), its nature depending on filling, local and non-local Hubbard interactions. Most recently, experiments on B-doped Sn/Si(111) reported the realization of chiral superconductivity Ming et al. (2023) with critical temperatures close to $T_{c}\sim 10$K.
@@
The aim of this paper is to revisit the band structures of the materials Sn/Si(111), Pb/Si(111) and Sn/SiC(0001) in their $\sqrt{3}\times\sqrt{3}$-reconstructued surface phase with a focus on the strength of spin-orbit coupling (SOC) in Sec. II; we complement our investigation with Pb/SiC(0001) whose $\sqrt{3}\times\sqrt{3}$-phase has not been reported so far. We do not investigate many-body ground states involving magnetic or charge order using DFT, but rather study and analyze the non-interacting metallic bandstructures. We further analyze phonon spectra of these materials in Sec. III in order to evaluate the electron-phonon coupling (EPC) strength $\lambda$ in Sec. IV A relativistic study of EPC for Pb/Si(111) is presented in Sec. V. The discussion of our results is presented in Sec. VI, before the end with an outlook in Sec. VII. In Appendix C we present analogous results of the electronic and phononic properties as well as EPC for Sn/Ge(111) and Pb/Ge(111).
@@ II Results I: Electronic structure
The electronic band structure of the four materials with the exception of Pb/SiC(0001) has been reported previously Tresca et al. (2018); Glass et al. (2015); Badrtdinov et al. (2016, 2018); Hansmann et al. (2013); Adler et al. (2019). Here we revisit and compare the four materials. We are interested in their metallic, non-interacting surface bandstructures. By deriving precise tight-binding models of these bands, we provide valuable input for any future many-body analysis of magnetic, charge-ordered, superconducting or otherwise ordered ground states. While the DFT bandstructures of the four materials are rather similar, it turns out that they realize different regimes of SOC ranging from 10% to 45% of the nearest-neighbor hopping amplitude $t_{1}$. This might have major implications for the (strongly) correlated many-body ground states of these materials, as pointed out in a recent study for the plain-vanilla triangular-lattice Rashba-Hubbard model Bunney et al. (2024).
@@
Table 2: Rashba SOC parameters using DFT+$U$. Parameters are given as ratio of $\alpha_{n}/t_{1}$ for $n$th nearest neighbor splitting.
@@
We use the ab initio package VASP Kresse and Furthmüller (1996) within the Perdew-Burke-Ernzerhof (PBE) generalized gradient approximation (GGA) Perdew et al. (1996) to compute the relativistic electronic bandstructures of the four materials. In particular, we employ DFT$+U$ using the approach of Dudarev et. al. Dudarev et al. (1998) where the Coulomb potential $U$ is applied only to the Si or SiC substrate atoms, but not on the Sn and Pb adatoms. The main effect of $U$ is to vary the gap size of the semiconductor substrate; only through changes of the gap size is the metallic surface band influenced. However, the surface band remains sufficiently uncorrelated within our first-principles calculations. The values for $U$ to get the correct band gaps are $U_{\rm Sn/Si(111)}=-2.84$ eV, $U_{\rm Pb/Si(111)}=-3.17$ eV, $U_{\rm Sn/SiC(0001)}=-11.64$ eV and $U_{\rm Pb/SiC(0001)}=-11.64$ eV.
DFT, particularly in its local density and gradient-corrected local density approximations, is known to inadequately describe the exact exchange-correlation hole. The exchange-correlation hole represents the probability distribution of finding an electron at position $r_{1}$ given the presence of another electron at position $r_{2}$. It is intuitive to consider that electronic correlations significantly influence the shape and profile of the exchange-correlation hole. By comparing numerically exact exchange-correlation holes with those obtained via DFT in simple systems, it becomes evident that the local density approximation underestimates the hole in regions near the atomic core, where states are more localized. To satisfy the probability sum rule, this underestimation results in a corresponding overestimation at greater distances, in regions dominated by delocalized states Gunnarsson et al. (1979). For strongly localized states, the DFT+$U$ method (with positive $U$) enhances the exchange-correlation hole by maintaining electrons at greater reciprocal distances. Conversely, in delocalized states, a correction can be achieved with a negative $U$ value, which effectively reduces electron-electron interactions. In small band-gap semiconductors like Si and SiC, where the band gap arises from delocalized $p$-like electronic states, reducing the effective Coulomb repulsion (via a negative $U$) increases the overlap of electron wavefunctions. This, in turn, modifies the bonding-antibonding interactions and adjusts the associated band gap. Alternative exchange-correlation functionals, such as hybrid functionals Heyd et al. (2003); Adamo and Barone (1999), address the local density approximation’s shortcomings by incorporating a fraction of exact exchange, which mitigates the self-interaction problem. However, the computational cost of exact exchange is prohibitive, especially for systems comprising 40 atoms with dense $k$-point sampling, such as adatoms on Si(111) or SiC(0001). Meta-GGA functionals Tao et al. (2003); Perdew et al. (2009); Becke and Johnson (2006); Tran and Blaha (2009) also improve band-gap estimates by incorporating the kinetic energy density into the exchange-correlation potential. Nonetheless, their application to large supercell models involving surfaces and vacuum regions is not yet standard practice. These aspects led us to opt for the accurate and widely adopted DFT+$U$ approach.
In these calculations we have used a plane-wave cutoff energy of 600 eV and energy convergence criteria of $10^{-8}$eV. For X/SiC(111) $7\times 7\times 1$ $\Gamma$-centered grids were used to sample the Brillouin zone, whereas for X/Si(111) $10\times 10\times 1$ $\Gamma$-centered grids were used for better tight-binding fits close to the $\Gamma$-point. Atomic positions within the cell were based off experimental values and then optimized within VASP. A force criteria of 0.0001 Å/eV was used in the relaxation procedures. For the X/SiC(0001) materials, the cell volume was also optimized to relieve excess pressure on the cell. In the relaxed cells, the adatoms sit above the bulk at heights of $h=2.48\,{\rm\AA}$ for Sn/Si(111), $h=2.62\,{\rm\AA}$ for Pb/Si(111), $h=2.10\,{\rm\AA}$ for Sn/SiC(0001), and $h=2.21\,{\rm\AA}$ for Pb/SiC(0001). Approximately 9.5 $\rm\AA$ and 2.6 $\rm\AA$ of vacuum were used for the X/Si(111) and X/SiC(0001) materials, respectively. The DFT$+U$ bandstructures along with the extracted tight-binding bands (discussed in detail below) are shown in Fig. 1, zoomed in on the energy range of the metallic surface band. We find the bandwidths $W/t_{1}=9.71$ ($W=0.51$eV) for Sn/Si(111), $W/t_{1}=7.69$ ($W=0.53$eV) for Pb/Si(111), $W/t_{1}=9.29$ ($W=0.40$eV) for Sn/SiC(0001) and $W/t_{1}=9.97$ ($W=0.40$eV) for Pb/SiC(0001). In Appendix A we show in Fig. 7 the same bandstructure plots with larger energy range involving substrate bands. The resulting Fermi surfaces are included as insets in Fig. 1. The Fermi surfaces of the X/Si(111) materials take a more warped hexagonal shape compared with the X/SiC(0001) materials due to the larger further nearest neighbor hopping values.
@@
Table 3: Radial Rashba SOC parameters present for 4th-, 7th-, 9th-nearest neighbor shells.
The four considered materials constitute a case where the metallic surface band is energetically completely separated from the substrate bands, and the description as a single-band tight-binding model is fully justified. The corresponding tight-binding parameters are listed in Tab. 1 and in Tab. 2. The tight-binding parameters are obtained by projecting Wannier functions onto the DFT bandstructures. Maximally localized Wannier functions are obtained using the Wannier90 softwarePizzi et al. (2020), initially projecting onto the adatom $p_{z}$ orbitals.
While the isolated triangular lattice of surface adatoms carries a $D_{6h}$ point group symmetry, the presence of the Si(111) or SiC(0001) substrate breaks inversion symmetry, and reduces the rotational symmetry from $C_{6}$ to $C_{3}$. This reduces the point group symmetry to $C_{3v}$. Group theoretic techniques were used to decompose the Hamiltonian terms, implementational details of which we delegate to Appendix A. Further symmetry analysis of the model has consequences discussed further below. Interestingly, we find for the nearest-neighbor Rashba amplitude $\alpha_{1}/t_{1}$ the value 0.10 for Sn/Si(111), 0.17 for Sn/SiC(0001), 0.29 for Pb/Si(111) and 0.44 for Pb/SiC(0001). The SOC splitting in Sn/SiC(0001) was previously reported to be $13.7\%$Glass et al. (2015). We attribute the increase in SOC strength in the X/SiC(0001) materials compared to their X/Si(111) counterparts due to the lack of inversion symmetry in the bulk SiC compared to bulk Si. The tight-binding Hamiltonian for our systems can be written as
@@
Table 4: Valley Zeeman parameters present for nearest neighbor shells with sites lying off the $C_{3v}$ mirror planes determined by the bulk substrate.
@@
We only find finite $\phi$ mixings for further-nearest neighbor shells containing 12 sites, where the radial strengths $\beta_{ij}$ shown in Tab. 3 are comparable to the conventional $\alpha_{ij}$. We also find what appears to be a valley Zeeman term, given by $H_{\rm{VZ}}$. This term is finite for nearest neighbors that do not lie on any of the mirror planes of the lattice. The sign of the term is governed by the form factor $\nu_{ij}=-{\rm{sgn}}(\sin(3\theta_{ij}))$. The strengths $J_{ij}$ given in Tab. 4 show that this effect is quite weak, with $J_{2}/t_{1}$ less than 1% in all materials.
@@
Figure 2: Top-down view of Sn/Si(111) with nearest neighbor Sn surface atoms (silver). Surface Si atoms (blue) have been darkened, emphasising how the bulk determines the choice of mirror planes (black lines) of the $C_{3v}$ symmetry.
@@
Some of our derived tight-binding parameter deviate significantly from previously reported ones Li et al. (2013); Adler et al. (2019); Badrtdinov et al. (2016, 2018); Glass et al. (2015); Hansmann et al. (2013). Most of the change is caused by the inclusion of $U$ in the DFT calculations. By increasing the gap size in order to match the experimentally observed values, the propagation of the Wannier functions is suppressed. The larger band gap acts as a wall resulting in highly localized Wannier functions and a decrease in hybridization with the bulk. Consequently our Wannier functions (shown in Fig. 3 a-d) are much more compact as compared with previous studies Badrtdinov et al. (2016); Glass et al. (2015). This leads to a reduction in the longer-ranged hoppings, e.g.,  $t_{2}/t_{1}$ from $-0.389$Adler et al. (2019) to $-0.270$ for Sn/Si(111), corresponding to a 30% decrease. While second and further neighbor hoppings are still non-negligible, these materials are closer to the nearest-neighbor triangular lattice than previously thought. More importantly, Rashba SOC can be quite sizeable for some of these materials. We note, however, that also our DFT calculations without inclusion of any $U$ led to some discrepancies with the literature. We assume that this is due to the higher resolution used in this work, as well as differing exchange-correlation functionals in the DFT calculations. In Appendix A we compare in Tab. 6 our derived parameters with previous works.
@@
As previously mentioned, the presence of the substrate reduces the point group symmetry of the surface triangular lattice from $D_{6h}$ to $C_{3v}$, by breaking inversion symmetry and reducing the $C_{6}$ rotation symmetry to a $C_{3}$. This can be seen in Fig. 2, which shows that the substrate also preserves the mirror planes which do not/do go through the bonds. We can further understand the terms in the Hamiltonian (1) in three subsets by a hierarchy of symmetry breaking. The kinetic Hamiltonian $H_{0}$ preserves the full $D_{6h}$ point symmetry group. The Rashba SOC terms $H_{R}$ only break inversion symmetry and therefore have a $C_{6v}$ point group symmetry. Note that the radial Rashba effect is only $C_{6v}$ symmetric on 12 site coordination shells. The valley Zeeman-type terms $H_{\rm{VZ}}$ additionally break the the rotational symmetry and have a $C_{3v}$ symmetry. Anisotropic valley Zeeman terms symmetric under $C_{3v}$ are only possible on nearest-neighbor shells where the mirror planes do not intersect with bonds, and hence are only finite on these shells. We also note that additional Rashba coupling terms would be allowed under the reduction to a $C_{3v}$ point group symmetry, but these were not observed.
@@
This symmetry hierarchy serves as one explanation of the relative strengths of the fitted parameters Winkler (2003). Initially employed to understand bandstructures of bulk semiconductors under external magnetic field or uniaxial strain Trebin et al. (1979); Suzuki and Hensel (1974), the symmetry hierarchy is the principle that terms in the Hamiltonian with lower symmetry typically have smaller prefactors than those with higher symmetry. While this would generally mean that the Hamiltonian can be well approximated by the higher symmetry terms alone, the lower symmetry terms may still lead to important physical effects. In this case, the Rashba and the anisotropic valley Zeeman type terms are small, but follow Moriya’s rules for antisymmetric anisotropic spin exchange Moriya (1960a). The Dzyaloshinskii–Moriya interaction Moriya (1960a, b); Dzyaloshinsky (1958) is an important factor for explaining weak anisotropic ferromagnetism in bulk materials, and is shown to drive a first order phase transition to a weakly ferromagnetic phase above a (second-order transition to) an antiferromagnetic phase. These types of interactions remain unexplored in the realm of surface states, and could have potential consequences for known spin-density wave phases in these materials Li et al. (2013).
@@
Figure 3: Maximally localized Wannier functions (MLWFs) for a Sn/Si(111), b Pb/Si(111), c Sn/SiC(0001) and d Pb/SiC(0001) are shown from top and side views. Adatoms are shown in grey, Sn in blue, C in brown, and H in pink. MLWFs are calculated using DFT+$U$ without the effects of SOC. We see low hybridization with the bulk in all materials, where the MLWFs only penetrate past the first layer.